Determine the slope and the y-intercept of the graph of the equation.

x + y - 10 = 0

Group of answer choices

m = -1; (0, -10)

m = 1; (0, 10)

m = -1; (0, 10)

m = 0; (0, 10)
Answer #1

Given x+y-10=0 y y = -x+10 componing with y = mate we have m= -1 and (= 10 Ans m=-) ; 10,10)